Question text Suppose that the average bill for the drive through of 24 customers at Burger Palace is $14.30 $14.30 . It is known that the population standard deviation is σ=2.30 σ=2.30 . Estimate the true mean bill for a customer at Burger Palace with 95% 95% confidence. Select one: a. $13.31<μ<$15.29 $13.31<μ<$15.29 b. $13.33<μ<$15.27 $13.33<μ<$15.27 c. $13.38<μ<$15.22 $13.38<μ<$15.22 d. $13.36<μ<$15.24
Answer #1
sample mean 'x̄= 14.300
sample size    n= 24.00
std deviation σ= 2.300
std error ='σx=σ/√n= 0.4695
for 95 % CI value of z= 1.960
margin of error E=z*std error = 0.920
lower bound=sample mean-E= 13.38
Upper bound=sample mean+E= 15.22

option C is correct

c. $13.38<μ<$15.22